[title Sol Summary – November 16th]
Crew 215 Sol Summary Report 16-NOV-2019
Summary Title: Data Consolidation
Author’s name: Andrew Wheeler
Mission Status: Active
Sol Activity Summary: The previous evening’s emergency drill has provided valuable
insights for the development of procedures applicable to operations in remote and
isolated locations. It was very satisfying to know we could carry out the tasks needed
to successfully complete the recovery. The morning allowed time to respond to questions
sent in from Australia, New Zealand and Papua New Guinea. It was a challenge. The Green
Hab continues to thrive with new herb and vegetable shoots appearing daily. After
lunch, Shane, Larissa and Guy returned to the micrometeorite sample grid in the stream
wash at MM5 (518462E 4248949N) where the remaining 36 samples were collected. We have
now sampled 100% of the grid and it has been retrieved. Dinner was a reheated rendition
of a Thai green curry from the previous evening and very welcome.
Look Ahead Plan: Tomorrow sees a day off for all crew with personal and hab chores to
be done. From Monday, Eva activities will wind back as data consolidation intensifies as
we move closer to rotation end.
Anomalies in work: N/A.
Weather: Marginally below freezing overnight, no wind, clear skies.
Crew Physical Status: Healthy
EVA: Micrometeorite sampling EVA. See EVA report
Reports to be file: Sol Summary Report, Operations Report, Science Report, GreenHab
Report, EVA Report, Journalist Report, Daily Photos.
Support Requested: N/A